# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 1.3. Morse theory with vertical boundary
For many constructions in the following sections, we will need a modified version of Morse theory to allow us to work with equivariant manifolds with boundary. The next two subsections are devoted to proving the foundational lemmas we will need for these constructions.
@@
Morse theory with vertical boundary uses Morse functions $f$ on manifolds $M$ with boundary where the associated gradient-like vector field of $f$ is everywhere tangent to $\partial M$. This should not be confused with the notion of a relative Morse function, for which $f$ is constant on connected components of $\partial M$ (e.g. on a cobordism). These two are not mutually exclusive, and one can have a relative Morse function with vertical boundary when $M$ is a manifold with corners. For instance, the height function $f$ on $D^{2}\times I$ given by $f(x,t)=t$ is such a function.
@@
It is important to note that we follow the convention of Kronheimer-Mrowka in [KM07] that the gradient-like vector field is given by $\nabla f$, rather than $\nabla f$. However, we do adapt some of the language from Borodzik-Némethi-Ranicki in [BNR16] to this convention.
@@
We say that an open submanifold $A$ of $M$ is fully vertical with respect to $f$ if for all $x\in A$, $\varphi(x,s)$ is also in $A$ for all $s$ in $\mathbb{R}$. In other words, $A$ is comprised of only trajectories, with no critical points, that are closed under the flow of $\nabla f$. Note that we will use “vertical” as a descriptor for various manifolds throughout the paper; this is not a technical term, and is mostly used in exposition to refer to manifolds that may have critical points, but to which $\xi$ is always tangent.
@@
We say that the index $\text{ind}(p)$ of a boundary critical point is $\dim\mathcal{D}(p)$. This definition agrees with the usual index of an interior critical point. However, note that with this definition, there can be no boundary stable critical points of index 0, nor boundary unstable critical points of index $n$.
@@
This definition should be interpreted as follows. Suppose $M$ is an $n$-manifold with boundary and fix a relative Morse function $f$ on $M$ such that $f$ is constant on some connected component of $\partial M$. Then an index $k$ right half-handle $R^{k}$ can be attached to that portion of $\partial M$ along $A^{k}$. However, we do this attachment so that $(\partial R)_{h}$ is horizontal with respect to $f$, and $(\partial R)_{v}$ is vertical with respect to $f$. Thus a right half-handle attachment changes both the topology of $\partial M$ and that of $M$ itself by attaching an index $k$ handle to each.
@@
Borodzik-Némethi-Ranicki showed, in analogy with classical Morse theory, that passing a boundary unstable critical point induces a right half-handle attachment. We will use this notion to build Seifert manifolds by hand by declaring that whenever we pass a critical point of the boundary, we attach a right half-handle to the Seifert manifold.
1.4. Equivariant Morse theory for finite group actions
We begin with a few remarks about equivariant Morse theory in general. For our purposes, and in much of the literature of equivariant Morse theory, the action of $\mathbb{Z}/d\mathbb{Z}$ on the codomain $\mathbb{R}$ is assumed to be trivial. Thus when we say that a Morse function $f$ is equivariant, we mean that $f\circ\tau=f$ where $\tau$ is the action on the domain.
Equivariant Morse theory for compact manifolds was studied by Wasserman in [Was69], building on work of Bott [Bot54] and Palais [Pal63]. Bao and Lawson [BL25] recently gave an extension of this result that shows genericity of Morse-Smale functions among pairs of equivariant functions and equivariant metrics. We will use this result implicitly throughout this subsection. Wasserman also showed, for compact Lie groups $G$, that the correspondence between critical points of a Morse function and handle attachment carries over to the $G$-equivariant setting by defining a critical $G$-orbit. In the case of our finite order rotation actions, equivariant handles always appear as either a single handle that is fixed setwise by $\tau$, or as a $d$-tuple of handles that are permuted by $\tau$.
It should be noted that in general, equivariant differential topology is much more restrictive than the non-equivariant setting. Transversality results often fail equivariantly. However, with sufficient conditions on the group action, the fixed set, and the manifold acted on, many standard results in Morse theory can be recovered in the equivariant setting. For instance, in the context of equivariant 4-manifold topology, Meier and Scott were able to prove an equivariant version of the Laudenbach-Poénaru theorem [MS25a], which enabled them to study equivariant trisections and bridge trisected surfaces [MS25b]. We will not use these results explicitly, but we mention them due to their close relationship to our study of equivariant 2-knots.
@@ 1.4. Equivariant Morse theory for finite group actions
Let $M^{k}$ be smoothly embedded in $B^{m}$ for $m-k\geq 2$ and equivariant with respect to the standard $\mathbb{Z}/d\mathbb{Z}$ action $\tau$ by rotation fixing a properly embedded copy of $B^{m-1}$. Let $f$ be an equivariant Morse function on $B^{m}$. For each critical point $p$ of $f$ restricted to $M$, let $\text{ind}(p)$ denote its index. We say that $M$ is in ordered equivariant Morse position with respect to $f$ if:
@@
    the critical points of $f|_{M}$ are isolated,
@@
    the Hessian of $f|_{M}$ at each critical point is nondegenerate,
@@
    every critical point of $f|_{M}$ is either fixed by $\tau$ or sent to another critical point of the same index, and
@@
    for all critical points $p_{i}$ and $p_{j}$, $f(p_{i})<f(p_{j})$ implies that $\text{ind}(p_{i})\leq\text{ind}(p_{j})$.
Recall that the purpose of our study of equivariant Morse theory is to build equivariant Seifert manifolds. In order to do this, we will need the boundary manifolds to be in ordered equivariant Morse position. For completeness, we prove a stronger statement that applies generally to embedded manifolds with finite order actions fixing a codimension 2 submanifold. This requires several steps, which are outlined in the following lemmas. The formal statement we wish to prove is:
Proposition 1.11.
Let $M$ be an $n$-manifold, possibly with boundary, embedded in $B^{m}$ where $m-n\geq 2$. Fix an equivariant Morse function $f$ on $B^{m}$ with two boundary critical points (one of index 0 and the other of index $m$), and let $\psi\colon M\to B^{m}$ be a $\tau$-equivariant embedding such that $A:=\operatorname{Fix}(\tau)\cap\psi(M)$ has at most one maximum and one minimum, with no other critical points with respect to $f$. Then $\psi$ is equivariantly isotopic to an embedding $\psi^{\prime}$ such that $\psi^{\prime}(M)$ is in ordered equivariant Morse position with respect to $f$, and $A^{\prime}:=\operatorname{Fix}(\tau)\cap\psi^{\prime}(M)$ has the same number and type of critical points as $A$.
The first step towards proving this is to describe when it is possible to perturb a Morse function on a manifold embedded in $\mathbb{R}^{n}$ to rearrange critical points in the complement of a fully vertical submanifold. This is essentially an adaptation of Theorem 4.1 of [BP16], and our proof follows theirs closely, but we here we are careful to keep the perturbation of the Morse function away from the fully vertical submanifold. We also only move one critical point of the Morse function restricted to the submanifold, deviating from [BP16] and other sources (e.g. [Mil65]) that contain this lemma that simultaneously rearrange a pair.
@@
Let $M$ be a compact $n$-manifold, possibly with boundary or corners, embedded in $W\cong B^{m-1}\times[0,1]$. Let $A$ be a properly embedded codimension 2 submanifold of $M$. Let $f\colon W\to[0,1]$ be a relative Morse function such that:
  1. (1)
    $f$ has no critical points on $W$ ($f$ is a submersion),
  2. (2)
    $f|_{M}$ has exactly two (possibly boundary) critical points, denoted $p_{k}$ and $p_{l}$ of indices $k$ and $l$ respectively, and
  3. (3)
    AA is fully vertical with respect to $f$.
Suppose that $k<l$ but $f(p_{k})>f(p_{l})$. Then there exists a Morse function $g$ on $W$ that also has no critical points on $W$, is homotopic to $f$ through submersions of $W$, and with $g(p_{k})<g(p_{l})$. Furthermore, $g$ can be chosen such that $f$ and $g$ agree on a neighborhood of $A$.
@@
For $p\in\{p_{k},p_{l}\}$, let $\mathcal{M}(p)=\mathcal{A}(p)\cup\{p\}\cup\mathcal{D}(p)$. Then since $k<l$, a codimension counting argument shows that $\mathcal{M}(p_{k})\cap\mathcal{M}(p_{l})$ is generically empty. Furthermore, since $A$ was assumed to be fully vertical, both $\mathcal{M}(p_{k})$ and $\mathcal{M}(p_{l})$ are disjoint from $A$.
Choose a neighborhood $N_{0}$ of $\mathcal{A}(p_{k})\cap(B^{m-1}\times\{0\})$ that is disjoint from a neighborhood of $\mathcal{A}(p_{l})\cap(B^{m-1}\times\{0\})$ and from a neighborhood of $A\cap(B^{m-1}\times\{0\})$. Let $\mu\colon B^{m-1}\times\{0\}\to\mathbb{R}$ be a smooth bump function supported on $N_{0}$ such that $\mu|_{\mathcal{A}(p_{k})\cap(B^{m-1}\times\{0\})}\equiv 1$. Then $\mu$ can be extended to a map $\hat{\mu}\colon W\to\mathbb{R}$ by extending along the flow of a gradient-like vector field of $f$. That is, if $\varphi\colon W\times\mathbb{R}\to W$ is the flow generated by $\nabla f$, then:
$\hat{\mu}(x)=\mu\left(\lim_{s\to-\infty}\varphi(x,s)\right)$
In particular, $\hat{\mu}$ is constant along trajectories of $f$. Let $N$ be the extension of $N_{0}$ to $W$ under this flow, so that $N$ is a neighborhood of $\mathcal{A}(p_{k})$ that is disjoint from neighborhoods of $\mathcal{A}(p_{l})$ and $A$.
Choose $a\in[0,1]$. This will be our target value for our new Morse function at $p_{k}$. We construct the new Morse function by rearranging the heights of points in $W$ based on their $\hat{\mu}-$value. This is done by choosing a function $\Upsilon\colon[0,1]\times[0,1]\to[0,1]$ satisfying:
@@
It is not enough to find new ambient Morse functions that rearrange critical points, however. For Proposition 1.11, we need these perturbations to be realized by an isotopy of the embedded manifold. We thank Mark Powell for his help with the following lemma.
@@
Let $M$ be a compact $n$-manifold, $W$ be a compact $m$-dimensional cobordism, and let $\psi$ be a smooth embedding of $M$ in $W$. Suppose that $f$ and $g$ are two relative Morse functions on $W$ that each have no critical points on $W$ and are homotopic through submersions of $W$. Then there exists an isotopy $\Psi\colon M\times[0,1]\to W$ and an orientation-preserving diffeomorphism $h\colon\mathbb{R}\to\mathbb{R}$ such that $\Psi(x,0)=\psi(x)$ and $\Psi(x,1)$ satisfies:
@@
We can now prove the existence of equivariant isotopies to ordered equivariant Morse position, given a fixed set with only two critical points.
@@
Recall that we begin with an equivariant Morse function $f$ on $B^{m}$, and that $A=\psi(M)\cap\operatorname{Fix}(\tau)$. By assumption, $f|_{A}$ has at most two critical points; denote these $p_{+}$ and $p_{-}$ according to whether they are maxima or minima. If $p_{+}$ exists, choose an equivariant neighborhood of $p_{+}$ and perturb $\psi$ to isotope it “up” along a path in $\operatorname{Fix}(\tau)$ such that $f(p_{+})>f(x)$ for all $x\in M$, using a bump function to smooth the corners. Such a path exists because the fixed set of the standard rotation action on $B^{m}$ is a copy of $B^{m-2}$, and thus is connected. Similarly, if $p_{-}$ exists, perturb $\psi$ to isotope an equivariant neighborhood of $p_{-}$ “down” so that $f(p_{-})<f(x)$ for all $x$ in $M$ and smooth the corners with a bump function. This procedure guarantees that the lowermost index 0 critical point and the uppermost index $n$ critical point of $M$ are both in $A$.
Let $\overline{M}$ be the quotient of $M$ by the pullback of $\tau$, and consider the quotient embedding $\overline{\psi}\colon\overline{M}\to\overline{B^{m}}$. (Note that $\overline{B^{m}}$ is still diffeomorphic to $B^{m}$; we use the symbol $\overline{B^{m}}$ to emphasize that we are working in the quotient manifold.) Since $f$ is equivariant, the image of $p_{+}$ and $p_{-}$ under the quotient map retain their maximal and minimal positions with respect to the induced Morse function $\overline{f}$ on $\overline{B^{m}}$.
@@
Let $a_{\pm}=\overline{f}(\overline{p}_{\pm})$, and let $\overline{W}=\overline{f}^{-1}([a_{-}+\epsilon,a_{+}-\epsilon])$ for $\epsilon$ small enough that the only two critical points of $\overline{f}|_{\overline{\psi}(\overline{M})}$ outside of $\overline{W}$ are $a_{+}$ and $a_{-}$. Then $\overline{M}\cap\overline{W}$ is a cobordism between two unknotted spheres (or disks, if $\partial M\neq\emptyset$) embedded in $B^{m-1}\times[0,1]$. The intersection of $\overline{\psi}(\overline{M})\cap W$ with $\operatorname{Branch}(\tau)$ is a copy of $S^{n-3}\times I$ embedded vertically with respect to $\overline{f}$. Denote this intersection $\overline{A}$, and note that because $\overline{f}$ is equivariant, $\overline{A}$ is fully vertical with respect to $\overline{f}$.
@@
Since $f$ has no critical points on $W$, we may use Lemma 1.12 to successively perturb $f$ to obtain Morse functions $f_{i}$ that rearrange pairs of critical points in the complement of $\operatorname{Branch}(\tau)$. The perturbations are supported on neighborhoods of descending manifolds of lower index critical points, and each $f_{i}$ is related to $f_{i+1}$ by a homotopy through submersions. Hence by Lemma 1.13, we may find an isotopy $\overline{\Psi}\colon\overline{M}\times[0,1]\to\overline{W}$ realizing this rearrangement, with $\overline{\Psi}(-,0)=\psi$. This can then be extended to an ambient isotopy (which, by abuse of notation, we also call $\overline{\Psi}$) by the Isotopy Extension Theorem. As we noted in Remark 1.15, since each perturbation of the Morse function was disjoint from $\operatorname{Branch}(\tau)$, $\overline{\Psi}$ can be taken to be supported in the complement of a neighborhood of $\operatorname{Branch}(\tau)$.
It remains to lift these rearrangements back to the cover. For this, we use Schwarz’s resolution of the Isotopy Lifting Conjecture [Sch80], which implies that isotopies in the quotient of a manifold by a compact Lie group can be lifted to isotopies of the manifold itself if they are supported away from the image of the fixed set. Thus we may lift $\overline{\Psi}$ to an isotopy $\Psi\colon M\times[0,1]\to B^{m}$ such that $\Psi(-,0)=\psi$ and the image $\Psi(M,1)$ is in ordered equivariant Morse position, again extended to an ambient isotopy. ∎
2. Equivariant Seifert solids
We now proceed to our construction of equivariant Seifert 3-manifolds. Our proof adapts a technique of Yanagawa [Yan69].

### v2

@@ 3. Embedded Morse theory and codimension 2 intersections
For many constructions in the following section, we will need a modified version of Morse theory to allow us to work with $n$-manifolds with boundary embedded in a cylinder of dimension $n+2$, while keeping track of a codimension 2 submanifold. This section is devoted to proving the foundational lemmas we will need for these constructions.
3.1. Review of Morse theory with vertical boundary
@@
Morse theory with vertical boundary uses Morse functions $f$ on manifolds with boundary $M$ together with gradient-like vector fields $\xi$ which are everywhere tangent to $\partial M$. This should not be confused with the notion of a relative Morse function, for which $f$ is constant on connected components of $\partial M$ (e.g. on a cobordism). These two are not mutually exclusive, and one can have a relative Morse function with vertical boundary when $M$ is a manifold with corners. For instance, the height function $f$ on $D^{2}\times I$ given by $f(x,t)=t$ is such a function.
@@
It is important to note that we follow the convention of Kronheimer-Mrowka in [KM07] that the gradient-like vector field mimics the behavior of $\nabla f$, rather than $\nabla f$. However, we do adapt some of the language from Borodzik-Némethi-Ranicki in [BNR16] to this convention.
@@
We introduce one more nonstandard definition which will be helpful for later sections.
Definition 3.1.
We say that an open submanifold $A$ of $M$ is fully vertical with respect to $f$ if for all $x\in A$, $\varphi(x,s)$ is also in $A$ for all $s$ in $\mathbb{R}$.
In other words, a fully vertical submanifold is closed under the flow of $\nabla f$. Note that we will use “vertical” as a descriptor for various manifolds throughout the paper; this is not a technical term, and is mostly used in exposition to refer to manifolds that may have critical points, but to which $\xi$ is always tangent.
@@
We say that the index $\operatorname{ind}(p)$ of a boundary critical point is $\dim\mathcal{D}(p)$. This definition agrees with the usual index of an interior critical point. However, note that with this definition, there can be no boundary stable critical points of index 0, nor boundary unstable critical points of index $n$.
@@
This definition should be interpreted as follows. Suppose $M$ is an $n$-manifold with boundary and fix a relative Morse function $f$ on $M$ such that $f$ is constant on some connected component of $\partial M$. Then an index $k$ right half-handle $R^{k}$ can be attached to that portion of $\partial M$ along $A^{k}$. However, we do this attachment so that $(\partial R)_{h}$ is horizontal with respect to $f$, and $(\partial R)_{v}$ is vertical with respect to $f$. Thus a right half-handle attachment changes both the topology of $\partial M$ and that of $M$ itself by attaching an index $k$ handle to each. An example right half-handle attachment is shown in Figure 8.
@@
Borodzik-Némethi-Ranicki showed, in analogy with classical Morse theory, that passing a boundary unstable critical point induces a right half-handle attachment. We will use this notion to build Seifert manifolds by hand in Section 4 by declaring that whenever we pass a critical point of the boundary, we attach a right half-handle to the Seifert manifold.
3.2. Morse theory with codimension two intersections
The goal of this subsection is to extend some of the standard results of embedded Morse theory to a setting in which we must keep track of a codimension 2 submanifold. Throughout, we fix the following setup:
  • $M$ is a smooth compact $n$-manifold, possibly with boundary or corners,
  • $W$ is a copy of $B^{n+1}\times I$, thought of as a manifold with corners,
  • $\psi$ is a smooth embedding of $M$ into $W$ such that there exists a decomposition of $\partial M$ into connected components and components joined by corners such that each component is either fully contained in $\partial W$, fully disjoint from $\partial W$ and has no corners, or intersects $\partial W$ only in its corners under $\psi$,
  • $Z$ is the subset $B^{n-1}\times I$ in $W$,
  • $A\subset M$ is a properly embedded copy of either $B^{n-2}$ or $S^{n-2}$ such that $\psi(M)\cap Z=\psi(A)$, and
  • $f\colon W\to I$ is a relative Morse function on $W$ such that $f$ has no critical points on $W$, $f|_{Z}$ has no critical points, $f|_{\psi(M)}$ is Morse-Smale, and $A$ is fully vertical with respect to $f|_{\psi(M)}$ away from at most one maximum and one minimum.
[…]
Figure 9. Schematic for the setup for Subsection 3.2 (in the case that $M$ has boundary and corners), with $f$ being projection onto the vertical axis. Note the flow lines of $\nabla(f|_{\psi(M)})$ showing that $A$ is fully vertical with respect to $f|_{\psi(M)}$. Note that this figure is not to scale: $\psi(M)$ and $Z$ are both codimension 2 in $W$, and $A$ is codimension 2 in $M$.
This is a more restricted setting than appears in the literature, and so we must reprove several standard facts from standard Morse theory. Recall the following standard definition.
@@ 3.2. Morse theory with codimension two intersections
We say that $\psi(M)$ is in ordered Morse position with respect to $f$ if:
@@
    the critical points of $f|_{\psi(M)}$ are isolated,
@@
    the Hessian of $f|_{\psi(M)}$ at each critical point is nondegenerate, and
@@
    for all critical points $p_{i}$ and $p_{j}$ of $f|_{\psi(M)}$, $f(p_{i})<f(p_{j})$ implies that $\operatorname{ind}(p_{i})\leq\operatorname{ind}(p_{j})$.
Given our setup above, we want to be able to isotope the map $\psi$ to an embedding where $\psi(M)$ is in ordered Morse position with respect to $f$. However, we also want this isotopy to fix a neighborhood of $A$, so that the image of $A$ is still fully vertical under the final map. Using the convention for isotopies that $\Psi_{t}(x):=\Psi(x,t)$, the formal statement we want to prove is:
Proposition 3.5.
There exists an embedding $\psi^{\prime}\colon M\to W$ and an isotopy $\Psi\colon M\times I\to W$ between $\psi$ and $\psi^{\prime}$ such that:
@@
    $\psi^{\prime}(M)$ is in ordered Morse position with respect to $f$,
  2. (2)
    $f|_{\psi^{\prime}(M)}$ has the same number and type of critical points as $f|_{\psi(M)}$, and
  3. (3)
    $\Psi_{t}\equiv\psi$ for all $t\in I$ on a neighborhood of $A$.
We prove this in two steps. First, we show that it is possible to perturb the standard Morse function $f$ on $W$ to another Morse function which, when restricted to $M$, has ordered critical points, while avoiding perturbation on a fully vertical codimension 2 submanifold of $M$. We then show that it is possible to achieve this same rearrangement via isotopy of $M$ inside $W$, so that $M$ is in ordered Morse position with respect to the original Morse function $f$. Because the perturbation of the Morse function in the first step is supported outside a neighborhood of $A$, the isotopy we construct in the second step fixes $A$ pointwise, and so $A$ remains fully vertical with respect to the restriction of $f$ to the image of $M$ throughout the isotopy.
Below we have the first step. This is essentially an adaptation of Theorem 4.1 of [BP16], and our proof follows theirs closely, but we here we are careful to keep the perturbation of the Morse function away from the fully vertical submanifold. We also only move one critical point of the Morse function restricted to the submanifold, deviating from [BP16] and other sources (e.g. [MIL65]) that contain this lemma that simultaneously rearrange a pair of critical points. We use the notation for homotopies that $F_{t}(x):=F(x,t)$.
@@
Suppose that $f|_{\psi(M)}$ has critical points $p_{k}$, $p_{j}$ such that $k<l$ but $f(p_{k})>f(p_{l})$. Then there exists a Morse function $g$ on $W$ and a homotopy $F\colon W\times I\to\mathbb{R}$ such that $F_{0}=f$ and $F_{1}=g$, $F_{t}$ has no critical points on $W$ for all $t$, $g(p_{k})<g(p_{l})$, and with $F_{t}\equiv\psi$ for all $t$ on a neighborhood of $A$.
@@
For $p\in\{p_{k},p_{l}\}$, let $\mathcal{M}(p)=\mathcal{A}(p)\cup\{p\}\cup\mathcal{D}(p)$. Then since $k<l$, a codimension counting argument shows that $\mathcal{M}(p_{k})\cap\mathcal{M}(p_{l})$ is generically empty. Furthermore, since $A$ was assumed to be fully vertical, both $\mathcal{M}(p_{k})$ and $\mathcal{M}(p_{l})$ are disjoint from $A$, as each of these submanifolds are fully vertical with respect to $f$ and thus closed under the flow of $\nabla f$.
Choose a neighborhood $N_{0}$ of $\mathcal{A}(p_{k})\cap(B^{n-1}\times\{0\})$ that is disjoint from a neighborhood of $\mathcal{A}(p_{l})\cap(B^{m-1}\times\{0\})$ and from a neighborhood of $A\cap(B^{m-1}\times\{0\})$. Let $\mu\colon B^{m-1}\times\{0\}\to\mathbb{R}$ be a smooth bump function supported on $N_{0}$ such that $\mu|_{\mathcal{A}(p_{k})\cap(B^{m-1}\times\{0\})}\equiv 1$. Then $\mu$ can be extended to a map $\nu\colon W\to\mathbb{R}$ by extending along the flow of a gradient-like vector field of $f$. That is, if $\varphi\colon W\times\mathbb{R}\to W$ is the flow generated by $\nabla f$, then:
$\nu(x)=\mu\left(\lim_{s\to-\infty}\varphi(x,s)\right)$
In particular, $\nu$ is constant along trajectories of $f$. Let $N$ be the extension of $N_{0}$ to $W$ under this flow, so that $N$ is a neighborhood of $\mathcal{A}(p_{k})$ that is disjoint from neighborhoods of $\mathcal{A}(p_{l})$ and $A$.
Choose $a\in[0,1]$. This will be our target value for our new Morse function at $p_{k}$. We construct the new Morse function by rearranging the heights of points in $W$ based on their $\nu$-value. This is done by choosing a function $\Upsilon\colon[0,1]\times[0,1]\to[0,1]$ satisfying:
@@
We thank Mark Powell for his help with the following lemma, which details the second step of the outlined procedure.
@@
Suppose that $f$ and $g$ are two relative Morse functions on $W$ that each have no critical points on $W$ and are homotopic through submersions of $W$. Then there exists an isotopy $\Psi\colon M\times[0,1]\to W$ and an orientation-preserving diffeomorphism $h\colon\mathbb{R}\to\mathbb{R}$ such that $\Psi(x,0)=\psi(x)$ and $\Psi(x,1)$ satisfies:
@@
We can now prove the existence of isotopies to ordered Morse position in the complement of a codimension two intersection.
@@
Recall that we begin with a Morse function $f$ on $W$, and that $\psi(A)=\psi(M)\cap Z$. By assumption, $f|_{\psi(A)}$ has at most two critical points; denote these $p_{+}$ and $p_{-}$ according to whether they are maxima or minima. If $p_{+}$ exists, choose a neighborhood of $p_{+}$ and perturb $\psi$ to isotope it vertically along a path tangent to $\nabla f$ in $Z$ such that $f(p_{+})>f(x)$ for all $x\in\psi(M)$, using a bump function to smooth the corners. Similarly, if $p_{-}$ exists, perturb $\psi$ to isotope a neighborhood of $p_{-}$ vertically down along a path tangent to $\nabla f$ so that $f(p_{-})<f(x)$ for all $x$ in $\psi(M)$ and smooth the corners with a bump function. This procedure guarantees that the lowermost index 0 critical point and the uppermost index $n$ critical point of $f|_{\psi(M)}$ are both in $\psi(A)$.
@@
Let $a_{\pm}=f(p_{\pm})$, and let $a^{\prime}_{\pm}=a_{\pm}\mp\epsilon$ for $\epsilon$ small enough that the only two critical points of $f|_{\psi(M)}$ outside of $V:=[a_{-}^{\prime},a_{+}^{\prime}]$ are $a_{-}$ and $a_{+}$. Then $M\cap V$ is a cobordism between two unknotted spheres (or possibly disks, if $\partial M\neq\emptyset$) embedded in $B^{n+1}\times[a_{-}^{\prime},a_{+}^{\prime}]$.
@@
Since $f$ has no critical points on $W$, we may use Lemma 3.6 to successively perturb $f$ to obtain Morse functions $f_{i}$ that rearrange pairs of critical points in the complement of $A$. The perturbations are supported on neighborhoods of descending manifolds of lower index critical points, and each $f_{i}$ is related to $f_{i+1}$ by a homotopy through submersions. Hence by Lemma 3.7, we may find an isotopy $\Psi\colon M\times[0,1]\to W$ realizing this rearrangement, with $\Psi_{0}=\psi$. By the Isotopy Extension Theorem, this can then be extended to an ambient isotopy which, by abuse of notation, we also call $\Psi$ . As we noted in Remark 3.9, since each perturbation of the Morse function was disjoint from $Z$, $\Psi$ can be taken such that $\Psi_{t}\equiv\psi$ for all $t$ on a neighborhood of $A$. This completes the proof. ∎
3.3. The equivariant case
It is worth saying a few words about equivariant Morse theory for finite cyclic group actions, and detailing why we may content ourselves with the lemmas proved in the previous subsection when working with periodic 2-knots. It should be noted that in general, equivariant differential topology is much more restrictive than the non-equivariant setting. Transversality results often fail equivariantly. However, with sufficient conditions on the group action, the fixed set, and the manifold acted on, many standard results in Morse theory can be recovered in the equivariant setting. For instance, in the context of equivariant 4-manifold topology, Meier and Scott were able to prove an equivariant version of the Laudenbach-Poénaru theorem [MS25a], which enabled them to study equivariant trisections and bridge trisected surfaces [MS25b]. We will not use these results explicitly, but we mention them due to their close relationship to our study of equivariant 2-knots.
Firstly, note that for our purposes and in much of the literature of equivariant Morse theory, the action of $\mathbb{Z}/d\mathbb{Z}$ on the codomain $\mathbb{R}$ is assumed to be trivial. Thus when we say that a Morse function $f$ is equivariant, we mean that $f\circ\rho=f$ where $\rho$ is the action on the domain. This was the setting studied by by Wasserman in [WAS69], building on work of Bott [BOT54] and Palais [PAL63]. Also worth mentioning is work of Bao and Lawson [BL25] in which they gave an extension of Wasserman’s result that shows genericity of Morse-Smale functions among pairs of equivariant functions and equivariant metrics.
The key fact we take from Wasserman’s work is that for compact Lie groups $G$, the correspondence between critical points of a Morse function and handle attachment carries over to the $G$-equivariant setting by defining a critical $G$-orbit. In the case of our finite order rotation actions $\rho$ with $\operatorname{Fix}(\rho)$ either a sphere or a ball, equivariant handles always appear as either a single handle that is fixed setwise by $\rho$, or as a $d$-tuple of handles that are permuted by $\rho$. Therefore, we may work with handle decompositions of simply periodic manifolds simply by working with their quotients and then lifting to the cover, so long as those handle decompositions are well-behaved relative to the image of the fixed set. Furthermore, it is immediate that $\operatorname{Fix}(\rho)$ is fully vertical with respect to $f$ away from its maximum and minimum. This, combined with Schwarz’s theorem on lifting isotopies, allows us to port all the results from the previous section into the $\mathbb{Z}/d\mathbb{Z}$-equivariant setting. For instance, we may define and prove the existence of an ordered equivariant Morse position for simply periodic manifolds which is completely analogous to Definition 3.4. Again, we fix the following notation for this subsection:
  • $\tilde{M}$ is a smooth, compact, simply-$d$-periodic $n$-manifold, possibly with boundary or corners,
  • $\tilde{W}$ is a copy of $\tilde{B}^{n+1}\times I$, thought of as a manifold with corners, equipped with the linear order $d$ rotation $\rho$ fixing $\hat{B}^{n-1}\times I$,
  • $\tilde{\psi}$ is a smooth, equivariant embedding of $\tilde{M}$ into $\tilde{W}$ such that there exists a decomposition of $\partial\tilde{M}$ into connected components and components joined by corners such that each component is either fully contained in $\partial\tilde{W}$, fully disjoint from $\partial\tilde{W}$ and has no corners, or intersects $\partial\tilde{W}$ only in its corners under $\tilde{\psi}$,
[…]
  • $\tilde{f}\colon\tilde{W}\to I$ is a relative Morse function on $\tilde{W}$ such that $\tilde{f}$ has no critical points on $\tilde{W}$, $\tilde{f}|_{\operatorname{Fix}(\rho)}$ has no critical points, $\tilde{f}|_{\tilde{\psi}(\tilde{M})}$ is Morse-Smale, and $\hat{A}$ is fully vertical with respect to $\tilde{f}|_{\tilde{\psi}(\tilde{M})}$ away from at most one maximum and one minimum.
Definition 3.10.
We say that $\tilde{M}$ is in ordered equivariant Morse position with respect to $\tilde{f}$ if:
  1. (1)
    the critical points of $\tilde{f}|_{\tilde{M}}$ are isolated,
  2. (2)
    the Hessian of $\tilde{f}|_{\tilde{M}}$ at each critical point is nondegenerate,
  3. (3)
    every critical point of $\tilde{f}|_{\tilde{M}}$ is either fixed by $\rho$ or sent to another critical point of the same index, and
  4. (4)
    for all critical points $p_{i}$ and $p_{j}$ of $\tilde{f}$, $\tilde{f}(p_{i})<\tilde{f}(p_{j})$ implies that $\operatorname{ind}(p_{i})\leq\operatorname{ind}(p_{j})$.
The following is the equivariant counterpart to Proposition 3.5.